OpenStudy (anonymous):

Tracey uses one-fourth of a cup of brown sugar in each batch of cookies. If she makes thirteen batches of cookies, how much brown sugar will she use?

OpenStudy (texaschic101):

if she uses 1/4 for 1 batch, then for 13 batches, she would use 1/4 * 13. can you do the math for 1/4 * 13 ?

OpenStudy (texaschic101):

1/4 * 13 = 13/4 = 3 1/4 cups of brown sugar are needed for 13 batches of cookies